OpenStudy (anonymous):

The formula P = S(1/2)^(h/4795), is the atmospheric pressure, P, in mm of mercury at elevation "h" meters above sea level. If S is 760 mm, at what elevation is the pressure 42 mm?

OpenStudy (wolf1728):

P = 760 * (.5) ^ (h/4795) P/760 = (.5) ^ (h/4795) log (P/760) = (h / 4795) * log(.5) log (P/760) / log(.5) = (h/4795) log (42/760) / log(.5) = (h/4795) -1.2575643019 / -0.3010299957 = (h/4795) 4.1775381856 = (h/4795) h = 4.1775381856 * 4795 h = 20,031.2955997227 height = 20,031

OpenStudy (wolf1728):

According to a pressure calculator located here: http://www.altitude.org/air_pressure.php at 20,031 meters, the atmospheric pressure is 45 mm Hg so I think my answer is correct.